# Through the Lens of Human-Human Collaboration: A Configurable Research Platform for Exploring Human-Agent Collaboration

### v1

@@ 4.1.1. Experimental Protocol
Our case study adheres to the same experiment rules as the original one. The core task in the Shape Factory experiment for each participant is to maximize their wealth by producing and trading shapes to complete shape orders. At the start of a session, each participant is assigned a ”specialty” shape that they can produce at a lower cost compared with other shapes. Each participant is given an order list of shapes other than their specialty to complete and earn revenue. Participants can perform two primary actions: produce shapes in their factory and trade shapes using money with other participants. Each participant has access to a computer interface that allows them to send text messages and trade offers to other participants.
The central tension of the experiment lies in the designed interdependence, where their shape orders do not consist of their own specialty shape, and there is a limit on the number of shapes they can produce in each round. As a result, participants must balance cooperation (to acquire the shapes they need) with competition (to negotiate favorable prices). Because each specialty shape is assigned to more than one participant, they have choices about their trading partners, which supports dynamic market behaviors and strategic planning.
@@ 4.1.2. Re-Implementation of the Shape Factory Experiment
In our study, we downsize the original 10-participant setup to six participants: one human and five LLM agents, where the agents can be viewed as analogous to remote human collaborators. The down-sampling of participant size requires changes to several interrelated experiment parameters, including the shape production limits and the number of shape types, but we preserved the core strategic tensions.
The experiment was implemented using our platform’s core features. We leverage Experiment Configuration Language to create experiment-specific objects, such as Shape, with attributes of Type, Regular Cost, Specialty Cost, Time Cost, and Production Status, and to extend existing objects, such as Participant, with attributes like Specialty Shape, Inventory, and Order. We also used ECL to configure the participant interface, where we enabled all five visual modules and customized the content to align with the experiment’s terminology. The controller automatically handles shape production and trade activities by subtracting the money from the corresponding participants and putting the shape back into their inventory once production is completed.
@@
We implemented the five agents using GPT-4o with a single end-to-end thinking inquiry, and integrated them using the Agent Context Protocol with the experiment controller. The ACP delivers structured state summaries to each agent at a fixed time interval to trigger its thinking query. We decided on 15 seconds as a reasonably balanced interval through pilot testing. In the state summary, the experiment controller delivers both publicly available information, such as time and participant list, and the private states available to the particular agents, such as messages and trade offers.
The agent’s behavior is guided by a structured prompt instruction, presented in Appendix D, that was iteratively refined through rigorous testing. This prompt includes a persona profile, experiment rules, valid action schemas, and instructions to encourage more human-like interaction patterns. Through this process, we addressed abnormal behaviors observed in preliminary tests, such as overly formal language, repetitive messaging, and a lack of contextual memory in conversations, which ensured the agents’ behaviors are reasonable and believable for the study.
@@ 4.1.3. Case Study Setup
Table 1. Demographic information of participants’ in the case study.
@@
This study examined the effect of workspace awareness. In the Control condition, participants had access to the real-time information dashboard. In the Experimental condition, this module was disabled. As suggested by the workspace awareness theory (Gutwin and Greenberg, 2002), we hypothesize that the experimental group will lead to less efficient coordination, lower task performance, and lower perceived awareness.
@@
We employed a crossed, between-subject study design with Communication Level and Awareness Level as the two factors. 16 participants were recruited from social media and professional networks. Each participant is asked to participate in both case studies and randomly assigned to one condition in each case study. Each case study consists of two consecutive sessions.
@@
We pre-configured each session with a unique session ID, then provided the instructions to each participant and asked them to complete all four sessions in the provided order; they are also required to fill out a post-study survey every two sessions, and a final survey with respect to agents’ behaviors after completing all four sessions.
@@ 4.1.5. Findings
Our findings demonstrate the platform’s effectiveness in supporting controlled experiments and capturing significant differences in both behavioral outcomes and subjective perceptions.
The Platform Captures Significant Differences in Collaboration Outcomes
Our case studies on manipulating interaction controls revealed differential effects on participants’ collaboration outcomes, as well as consistent behavioral patterns across both conditions. In $CS_{CL}$, where the private messaging function was available only to the control group, disabling chat improved participants’ final wealth. As shown in the first chart in Figure 6, human participants in the Experimental group achieved higher final wealth (\$406.4, SD=66.7) than those in the Control group (\$365.8, SD=92.9). This result is accompanied by more successful trades in the Experimental group, suggesting that removing communication overhead allowed participants to remain focused on the task and yield more favorable outcomes. In contrast, in $CS_{AL}$, manipulating real-time awareness of other participants’ status had a less pronounced effect on human participants. Final wealth among human participants was nearly identical between the Control (\$375.1, SD=118.3) and Experimental (\$376.6, SD=123.0) groups, despite the Experimental group engaging in more trades and communication. Interestingly, this increase in activity did not translate into higher profitability. However, agent performance was negatively impacted in the Experimental group, with average wealth decreasing from \$254.0 to \$218.1. This result suggests that LLM agents are less able to strategically coordinate with other participants compared to humans.
Beyond these case-specific effects, we observed consistent behavioral patterns across both studies. First, as shown in the fourth charts in Figure 6 and 7, human participants’ final wealth steadily increased from the first to the fourth session, with a regression model confirming session order as a significant predictor (coef=+26.11, p=0.015). Second, correlation analysis revealed a positive link between final wealth and both the number of successful trades and trade efficiency ($r\approx 0.53$), and a weaker correlation with average trade price ($r\approx 0.33$). Participants’ communication behaviors showed mixed effects: while shorter response latencies and longer messages correlated with more successful trades, overall message volume negatively correlated with wealth. These findings demonstrate that our platform effectively supports longitudinal study designs and can capture the evolution of participant strategies and broader behavioral trends as they adapt to human-LLM-agent collaboration dynamics.
@@
Table 2. Paired t-test results for participants’ perceptions of distributed collaboration. Green arrows ($\uparrow$) indicate mean values above the midpoint, while red arrows ($\downarrow$) indicate those below. Statistically significant $p$-values are bolded, where * denotes $(p\leq 0.05)$, and ** denotes $(p\leq 0.01)$.
The Platform Supports Validated Measures of User Perception.
The post-study survey results, summarized in Table 2, reveal that participants’ perceptions of distributed collaboration were strongly associated with the level of support provided by the platform. In control groups, where communication and awareness were more supported, participants consistently reported higher levels of trust in their collaborators, greater awareness of others’ actions, and more positive evaluations of team effectiveness and interaction. These results suggest that our platform enables participants to build trust, maintain sufficient workspace awareness, and sustain effective teamwork. In terms of workload, participants reported significantly lower physical demand and frustration than the neutral point, allowing them to focus on the collaborative task rather than the interface itself. The platform’s usability was further supported by the performance ratings, which were consistently above the neutral point in both groups. In contrast, participants’ perceptions of the LLM agents were notably less favorable. Anthropomorphism was perceived negatively, with participants rating the agents as more machine-like (M=2.35, p¡0.05) and less lifelike (M=2.15, p¡0.01). Similarly, the perceived intelligence was also below neutral, with agents rated as more incompetent than competent (M=2.45, p¡0.05). We believe the lower ratings are reasonable as the agents were deliberately configured with only basic capabilities, and more advanced prompt-engineering techniques and richer persona design can be explored in future work.
@@ 4.2. Cognitive Walkthrough with HCI Researchers
The second part of our evaluation employed a participatory cognitive walkthrough to assess the usability of the researcher interface. The study was designed to identify potential workflow breakdowns and gather qualitative feedback from target users, who are HCI researchers with research interests in human-AI collaboration, to iteratively refine the interface for study setup and analysis.
@@ 4.2.1. Setup
We recruited five HCI researchers with expertise in human-AI collaboration through professional networks and snowball sampling (Goodman, 1961). Participants were all doctoral students, postdoctoral researchers, or research scientist in HCI with prior experience in human-AI collaboration research. Each session was conducted remotely via Zoom, moderated by the first author, and lasted roughly 45 minutes. The study was video recorded upon acquiring participants’ consent at the beginning.
ID Gender Age Group Occupation
P1 Female 30-35 Clinical Assistant Professor
P2 Female 25-30 PhD Student
P3 Male 30-35 Applied Scientist
P4 Male 25-30 Postdoc
P5 Male 25-30 Senior PhD student
Table 3. Participant demographic information for the cognitive walkthrough. All participants are experienced HCI researchers with expertise in human-AI collaboration.
The session began with a brief introduction to the platform’s motivation and a demonstration of the participant interface to familiarize the researcher with the experimental context. We then provide an instruction document, along with the link to the researcher interface, to the participant, who is assigned an imaginary persona of an HCI researcher interested in using our platform for a research study on how communication channels affect human-agent collaboration with the Shape Factory experiment. Two tasks are presented to the participants, each associated with one user scenario: (1) set up a new session for a Shape Factory experiment and (2) analyze the results from a pre-populated, completed session. Participants shared their screen and were instructed to think aloud as they navigated the interface to complete these two scenarios. The facilitator observed, took notes on usability issues, and asked probing questions. We conducted a short semi-structured interview after the walkthrough to gather overall impressions and discuss potential research applications. All study procedures were approved by the first author’s Institutional Review Board.
The recordings of each session were transcribed and analyzed alongside the facilitator’s notes, where we conducted a thematic analysis to identify recurring usability challenges, comprehension difficulties, and design suggestions.
Preliminary Researcher Interface
The preliminary research interface was designed for configuring the case studies presented in Section 4.1; thus, all the necessary functionalities are implemented and available to the participants. We present the original interface design in Appendix E. The researcher interface organizes the experimental workflow into four tabs: experiment selection, session parameters, real-time monitoring, and post-study analysis. A short text description was presented along with all the experiment-specific parameters when the researchers chose one experiment in the first tab; then, they were directed to the session parameters page, which organizes the controls into two groups: interaction variables and participant registration. In the original workflow, users will create a session after they have specified all the parameters and controls, but before registering participants.

### v2

@@ 4.1.1. Experimental Protocol
Our case study adheres to the same experiment rules as the original one. The core task in the Shape Factory experiment for each participant is to maximize their wealth by producing and trading shapes to complete shape orders. At the start of a session, each participant is assigned a ”specialty” shape that they can produce at a lower cost compared with other shapes. Each participant is given an order list of shapes without their specialty to complete and earn revenue. Participants can perform two primary actions: produce shapes in their factory and trade shapes using money with other participants. Each participant has access to a computer interface that allows them to send text messages and trade offers to other participants.
The central tension of the experiment lies in the interdependence, where each participant’s shape orders do not consist of their own specialty shape, and there is a limit on the number of shapes they can produce. As a result, participants must balance cooperation (to acquire the shapes they need) with competition (to negotiate favorable prices). Because each specialty shape is assigned to more than one participant, they can choose trading partners, which supports dynamic market behaviors and strategic planning.
@@ 4.1.2. Re-Implementation of the Shape Factory Experiment
In our study, we downsize the original 10-participant setup to six participants: one human and five LLM agents, where the agents can be viewed as analogous to remote human collaborators. The down-sampling of participant size requires adjusting several interrelated experiment parameters (e.g., the shape production limits and the shape types), but we preserved the core strategic tensions.
The experiment was implemented using our platform’s core features. We leverage Experiment Configuration Language to create experiment-specific objects, such as Shape, with attributes of Type, Regular Cost, Specialty Cost, Time Cost, and Production Status, and to extend existing objects, such as Participant, with attributes like Specialty Shape, Inventory, and Order. We also used ECL to configure the participant interface by enabling all five visual modules and customizing their content to match the experiment terminology. The controller automatically manages shape production and trades by deducting money from the corresponding participants and returning completed shapes to their inventories.
@@
We implemented the five agents using GPT-4o with a single end-to-end thinking inquiry, and integrated them using the Agent Context Protocol with the experiment controller. The ACP delivers structured state summaries to each agent at a fixed time interval to trigger its thinking query. We decided on 15 seconds as a reasonably balanced interval through pilot testing. Each state summary includes both public information (e.g., time and participant list) and agent-specific private states (e.g., messages and trade offers).
The agent’s behavior is guided by a structured prompt (Appendix E) that was iteratively refined through rigorous testing. Through this process, we addressed abnormal behaviors observed in preliminary tests, such as overly formal language, repetitive messaging, and a lack of contextual memory in conversations, which ensured the agents’ behaviors are reasonable and believable for the study.
@@ 4.1.3. Case Study Setup
Table 2. Demographic information of case study participants.
Table 2. Demographic information of participants in the case study. This table summarizes the background information of participants, including their identifiers and roles in the study.
@@
This study examined how different levels of workspace awareness impact collaborative outcomes. In the Control condition, participants had access to the real-time information dashboard. In the Experimental condition, this module was disabled. As suggested by workspace awareness theory (Gutwin and Greenberg, 2002), we hypothesize that the experimental group will lead to less efficient coordination, lower performance, and lower perceived awareness.
@@
We employed a crossed, between-subject study design with Communication Level and Awareness Level as the two factors. 16 participants were recruited from social media and professional networks. Each participant participates in both case studies and is randomly assigned to one condition in each. Each case study consists of two consecutive sessions.
@@
We pre-configured each session with a unique session ID, then provided the instructions to each participant and asked them to complete all four sessions in the provided order; they are also required to fill out a post-study survey every two sessions, and a final survey with respect to their perception of agents’ behaviors after completing all four sessions.
@@ 4.1.5. Findings
Fig. 6. Collaboration outcomes in the CS_CL case study. A set of four bar charts shows session results in order from left to right: average wealth earned by participants, the number of successful trades completed, the total message volume exchanged, and the wealth distribution for human participants.
@@
Fig. 7. Collaboration outcomes in the CS_AL case study. Four bar charts illustrate outcomes across session order: participants’ average wealth, number of successful trades, message volume, and human participants’ final wealth.
Our analysis indicates the platform’s interaction controls isolate a communication variable and yield significant, measurable differences in both behavioral outcomes and subjective perception with the Shape Factory experiment.
In the Communication-Level study ($CS_{CL}$), when private chat was disabled (experimental group), human participants earned higher final wealth than when chat was enabled ($M_{control}=\mathdollar 365.8,\allowbreak SD=92.9;M_{experiment}=\mathdollar 406.4,SD=66.7$), with a mean difference of $40.6$. The comparison is shown in the first chart in Figure 6. Behavioral logs show a higher count of successful trades in the experimental group, suggesting that the platform effectively constrained the socio-emotional exchange through the communication channel and shifted participants’ behaviors from protracted negotiation to rapid, market-based signaling. The demonstrated trade-off between communication overhead and task focus opens up opportunities for researchers to analyze communication channel capacities in human-agent collaboration with our platform.
In contrast, the Awareness Level study ($CS_{AL}$) demonstrated the platform’s ability to reveal asymmetries between human and agent adaptability. The manipulation workspace awareness support through Information Dashboard produced very similar human outcomes across conditions ($M_{control}=\mathdollar 375.1,SD=118.3;\allowbreak M_{experimental}=\mathdollar 376.6,SD=123.0$). Agent outcomes, by contrast, received a considerable negative impact in the Experimental group, with average agent wealth decreasing from \$254.0 (SD=39.27) to \$218.1 (SD=38.80). This result critically demonstrated that the platform can identify differences between humans and agents about collaborative effectiveness. Specifically, reduced workspace awareness shared among collaborators hinders agent collaboration more than human collaboration, which motivates future research to conduct comparative evaluations of alternative agent architectures.
Furthermore, our platform demonstrated robustness for longitudinal analysis as we observed consistent outcome improvement over time across both studies. As shown in the fourth charts in Figure 6 and 7, human participants’ final wealth steadily increased from the first to the fourth session, with a regression model confirming session order as a significant predictor (coef=+26.11, p=0.015). Correlation analysis further revealed a positive link between final wealth and both the number of successful trades and trade efficiency ($r\approx 0.53$), and a weaker correlation with average trade price ($r\approx 0.33$). Together, these patterns indicate that the platform captures learning effects and strategic changes over time, which is essential for analyzing how humans adapt their strategies and alter their behavioral trends as they become more familiar and engaged in the human-agent collaboration dynamic.
@@
Table 3. Paired t-test results for participants’ self-reported perceptions in the Shape Factory experiments. Green arrows ($\uparrow$) indicate mean values above the midpoint, while red arrows ($\downarrow$) indicate those below. Statistically significant $p$-values are bolded, where * denotes $(p\leq 0.05)$, and ** denotes $(p\leq 0.01)$.
Table 3. Paired t-test results for participants’ perceptions of distributed collaboration in Shape Factory experiments. Columns list the measured dimension, mean, standard deviation, t value, p value, and order effect. Dimensions include perceived trust, workplace awareness, collaboration effectiveness, social presence, workload components (mental, physical, temporal demand, performance, effort, frustration), anthropomorphism, and perceived intelligence. Green arrows indicate mean values above the scale midpoint, and red arrows indicate values below it. Statistically significant p values are bolded (p ≤ .05 marked with *, p ≤ .01 with **).
The perception measures via self-reported survey, summarized in Table 3, also aligned with the hypothesis about communication and awareness manipulations. In control groups, where the platform enabled higher communication and awareness support, participants reported significantly higher levels of Perceived Trust ($p=0.03^{*}$) and Workspace Awareness ($p=0.01^{**}$). Usability and workload ratings indicated that the system itself did not impose undue burden, with Physical Demand and Frustration falling significantly below the neutral point across conditions ($p<0.01^{**}$).
Finally, our experiments provided a baseline for evaluating humans’ perceptions of LLM agents in collaborative tasks, where the participants rated the basic LLM agents as more ”machine-like” ($p<0.05$) and ”incompetent” ($p<0.05$). Such results are aligned with the basic, prompt-based agent design that does not explicitly target critical collaborative capacities in remote teams, such as maintaining common ground and awareness (Clark and Brennan, 1991; Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), repairing breakdowns (Schmidt and Bannon, 1992), and negotiating tasks (Bos et al., 2002, 2004). Taking together, these findings show that the platform’s controls produce expected directional changes in the human-agent collaboration dynamics and support both outcome-level and process-level in-depth analyses.
4.2. Case Studies with Hidden Profile Experiment
To demonstrate the platform’s capacity to support complex collaborative decision-making tasks involving information asymmetry, we conducted a between-subjects study with the Hidden Profile paradigm (Stasser and Titus, 1985, 1987). Unlike the Shape Factory, which focuses on resource negotiation, the Hidden Profile measures the team’s ability to pool distributed knowledge to reach an optimal decision.
4.2.1. Experimental Protocol.
The Hidden Profile task is a three-participant collaborative decision-making experiment. Each participant is provided with information about all decision alternatives, and one of them is superior to the others (i.e., has more advantages) when full information is available. However, the advantages of the best alternative are unshared, leading to the shared information supporting a suboptimal alternative. As a result, participants can identify the optimal alternative only after the unshared information is pooled through discussion and integrated into a revised decision.
We follow the core structure introduced by Schulz-Hardt and Mojzisch (2012), which asks each participant to role-play a personnel manager in an airline company that is looking to hire a new pilot for long-distance flights. Among all three candidates (denoted as A, B, and C), candidate C has seven advantages and three disadvantages, whereas candidates A and B each have four advantages and six disadvantages. In the experiment, the advantages of candidate C, the optimal candidate, are distributed and unshared among participants.
4.2.2. Re-Implementation of the Hidden Profile Experiment
To allow human participants to read the assigned candidate materials, we implemented a document-reading feature that enables in-interface preview of the assigned PDF files. Since the Hidden Profile experiment primarily requires participants to communicate and exchange information, we configured the participant interface using ECL by enabling the My Action module to display voting results, the My Task module to present the experiment rules, and setting the communication level to Group Chat. The interfaces for the Hidden Profile experiment are presented in Figure 15, 16, and 17. We implemented the agents using GPT-4o, integrated with ACP to allow agents to perceive the experiment state and execute actions.
4.2.3. Manipulation and Procedure.
We manipulated Agent Responsiveness, drawing on mixed-initiative interaction principles (Allen et al., 1999). In the passive condition, each agent only thinks and responds after the human participant sends a message. In the proactive condition, each agent automatically executed a perceive–predict cycle every 30 seconds to assess the conversational state, decide whether to engage in the conversation, and determine what message to send. The underlying policy and access to materials were identical across conditions. Prompt instructions, as shown in Appendix E, are iteratively tested and refined to eliminate faulty and obvious non-humanlike behaviors.
16 Participants are recruited through Amazon Mechanical Turk (MTurk)  [^4] https://www.mturk.com/ and randomly divided into two conditions (8 participants each). All participants first review the experiment rules and procedures, then read the assigned candidate materials for three minutes. Afterward, the human participant and two agents submit an initial vote, followed by a 15-minute group discussion and a final vote.
4.2.4. Measures.
The primary outcome was decision accuracy at the team level, defined as selecting the target candidate after discussion. We also measured human participants’ message volume and lengths in both conditions. Hidden profile research links accuracy to disclosure and integration of unique evidence, so these outcomes index whether the responsiveness control shapes information pooling in the intended direction. We analyzed the self-reported surveys using the same method as reported in Section 4.1.4.
In addition, we captured information sharing by labeling each atomic fact as shared or unique and recording the first mention, the number of mentions, and whether the fact was integrated into an argument for or against a candidate. We analyzed the time to the first mention of each unique advantage and the proportion of unshared unique facts mentioned at least once. To characterize conversational dynamics, we computed the share of agent-initiated turns, response latency distributions, turn-taking equality, question-asking, and summarization rates. This follows established coding traditions for hidden profile discussions and separates initiative from simple message volume.
4.2.5. Findings.
Our analysis shows that the agent responsiveness manipulation led to significant behavioral pattern changes in participants’ decision-making outcomes, in-experiment behaviors, and subjective perceptions in the Hidden Profile experiment.
In the Proactive Condition, the team-level accuracy reached 16.7%. By contrast, team-level accuracy of participants in the Passive Condition only achieved 8.3%. Analysis of the discussion logs suggests that proactive agents facilitated meaningful pooling of the unshared information. For example, one agent noted that “Candidate C’s interpersonal skills could foster teamwork, but their egocentric nature might create challenges,” while another added that “Candidate C’s ability to create a positive atmosphere could really enhance teamwork.” These complementary perspectives prompted the human participant to suggest “I think we need to find out more (details about C).” This pattern indicates that agent-initiated discussion may better support information sharing.
We acknowledge that both the Proactive and Passive Conditions have relatively low overall team accuracy. This result is indeed aligned with the classic Hidden Profile study findings (Stasser and Titus, 1985, 1987), where humans often engage in biased discussion, as the initial individual preferences and group consensus are misled by the provided materials because the supportive information for Candidate C is unshared. Regarding behavioral patterns, participants in the Proactive Condition sent an average of 13.1 (SD = 2.6) messages with a mean length of 77.2 (SD = 36.2), whereas participants in the Passive Condition sent fewer (M = 10.75, SD = 3.8) but longer messages ($M_{length}$ = 95.3, SD = 46.7). This pattern suggests that when the agent is passive, humans need to communicate more information per message to progress the discussion, indicating higher communication effort than in the Proactive Condition.
According to the self-reported survey (see Table 9), participants’ perception measures were also consistent with the agent responsiveness manipulation. In the Proactive Condition, participants reported significantly higher Perceived Trust ($p=0.01^{**}$). Usability and workload ratings further indicate that the system was easy to use: Physical Demand was significantly lower than the neutral point ($p\leq 0.01^{**}$), and Performance was significantly higher than the neutral point ($p\leq 0.02^{*}$) across both conditions.
Collectively, the findings of the case study on Shape Factory and Hidden Profile experiments with different interaction controls demonstrate the platform’s research efficacy of isolating interaction variables to reveal their impact on collaborative processes and outcomes across diverse domains in human-agent collaboration.
@@ 4.3. Cognitive Walkthrough with HCI Researchers
The second part of our evaluation employed a participatory cognitive walkthrough to assess the platform’s usability and learning curve for its target users. The study focused on the researcher experience: specifically, whether HCI researchers with no prior exposure to the system could successfully configure a valid study and interpret its results without extensive training. Additional information is collected with respect to researchers’ potential research interests with the use of this system.
@@ 4.3.1. Setup
We recruited five HCI researchers through professional networks and snowball sampling (Goodman, 1961). Participants were all doctoral students, postdoctoral researchers, or research scientists in HCI with diverse levels of experience in human-AI collaboration research. Each session was conducted remotely via Zoom, moderated by the first author, and lasted roughly 45 minutes. The study was video recorded upon acquiring participants’ consent at the beginning, and all study procedures were approved by the first author’s Institutional Review Board.
ID Gender Age Group Occupation Year of Experience
P1 Female 30–35 Assistant Professor 5+ years
P2 Female 25–30 Junior PhD Student 3–5 years
P3 Male 30–35 Applied Scientist 5+ years
P4 Male 25–30 Postdoc 5+ years
P5 Male 25–30 Senior PhD Student 3–5 years
Table 4. Participant demographic information for the cognitive walkthrough. All participants are experienced HCI researchers with expertise in human–AI collaboration.
Table 4. Demographic information of participants in the cognitive walkthrough. All listed participants are experienced HCI researchers with expertise in human–AI collaboration.
The session began with a brief introduction to the platform’s motivation and a demonstration of the participant interface to familiarize the researcher with the experimental context. We assigned each participant an imaginary persona of an HCI researcher designing a study on communication channels in the Shape Factory paradigm with our system, along with the researcher interface link. Participants interact with a fully functional prototype of the Researcher Interface (see Appendix F for the original design). While this version contained all core functionalities used in the Case Studies, its information architecture differed slightly from the final design presented in Section 3.2.2. Two tasks are presented to the participants, each associated with one user scenario: (1) set up a new session for a Shape Factory experiment and (2) analyze the results from a pre-populated, completed session. Participants shared their screen and were instructed to ”think aloud” as they navigated the interface. The facilitator observed, took notes on usability issues, and asked probing questions. We conducted a short semi-structured interview after the walkthrough to gather overall impressions and discuss potential research applications.
The recordings of each session were transcribed and analyzed alongside the facilitator’s notes, where we conducted a thematic analysis to identify recurring usability challenges, comprehension difficulties, and design improvement opportunities.